OpenStudy (anonymous):

How do you evaluate this integral using integration by parts ? ∫In(x)^1/2

OpenStudy (anonymous):

to clarify is that (ln(x))^(1/2) or ln(x^(1/2))?

OpenStudy (anonymous):

i think its \[\int\limits \ln (\sqrt{x}) dx\]because there is no closed form for \[\int\limits \sqrt{\ln x} dx\]

OpenStudy (anonymous):

@mukushla yeah thats the question

OpenStudy (anonymous):

how would you solve it ?

OpenStudy (anonymous):

better to write it in the form\[\int\limits \ln (\sqrt{x}) dx=\frac{1}{2} \int\limits \ln x dx\] now for integration by parts let \(u=\ln x\) and \(dv=dx\)

OpenStudy (anonymous):

oh Yeaahh I got the rest . thanks alot :)

OpenStudy (anonymous):

very well........:)